OpenStudy (anonymous):

y^2=4x+4y , find vertex,latus rectum,directrix,focus of parabola

OpenStudy (accessdenied):

Ideally, we'd like our equation to be in the form: \(x = a(y - k)^2 + h\). Do you know how to get the equation into this form?

OpenStudy (anonymous):

no

OpenStudy (accessdenied):

Okay, so we basically want to get the y-values on one side and the x-values on the other. We'll subtract 4y from both sides. y^2 - 4y = 4x Then, we will complete the square for "y^2 - 4y." The term we add to create the perfect square is half of the -4, squared, so... y^2 - 4y + 4 - 4 = 4x (y - 2)^2 - 4 = 4x Then we divide by 4. (1/4)(y - 2)^2 - 1 = x Or, if we reverse it... x = (1/4)(y - 2)^2 - 1.

OpenStudy (accessdenied):

Then, the vertex is the point "(h,k)" from "x = a(y - [k])^2 + [h]" k = 2, h = -1 So, the vertext is V(-1, 2). Does that make sense?

OpenStudy (anonymous):

no

OpenStudy (accessdenied):

Well, I'm sorry I was unable to help you. I hope you are able to figure it out.